3x/2 – 2x/3 = 83
x/2 – 2x/3 = 8
 
3x/2 – 2x/3 = 83
x/2 – 2x/3 = 8
3x/2 – 2x/3 = 83
x/2 – 2x/3 = 8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3x/2 – 2x/3 = 8

(9x – 4x)/6 = 8

5x = 48

x = 48/5

Thanks